OpenStudy (anonymous):

the ratio of girls to boys at school is 7:4. if there are currently 189 girls at the school, then how many additional boys would be needed so that the resulting ratio of girls to boys is 3:2? Show answer in two different ways...

OpenStudy (anonymous):

7x+4x+189 11x=189 x=189/11 x=17.2 7x/4x=120/69 3x+2x=189 5x=189 x=189/5 x=37.8 3x/2x=113/75.6

OpenStudy (anonymous):

please explain...